### v1

@@ Abstract
Social media platforms systematically reward popularity but not authenticity, incentivizing users to strategically tailor their expression for attention. We develop a utilitarian framework addressing strategic expression in social media. Agents hold fixed heterogeneous authentic opinions and derive (i) utility gains from the popularity of their own posts—measured by likes received, and (ii) utility gains (losses) from exposure to content that aligns with (diverges from) their authentic opinion. Social media interaction acts as a state-dependent welfare amplifier: light topics generate Pareto improvements, whereas intense topics make everyone worse off in a polarized society (e.g., political debates during elections). Moreover, strategic expression amplifies social media polarization during polarized events while dampening it during unified events (e.g., national celebrations). Consequently, strategic distortions magnify welfare outcomes, expanding aggregate gains in light topics while exacerbating losses in intense, polarized ones. Counterintuitively, strategic agents often face a popularity trap: posting a more popular opinion is individually optimal, yet collective action by similar agents eliminates their authentic opinion from the platform, leaving them worse off than under the authentic-expression benchmark. Preference-based algorithms—widely used by platforms— or homophilic exposures discipline popularity-driven behavior, narrowing the popularity trap region and limiting its welfare effects. Our framework fills a critical gap in the social media literature by providing a microfoundation for user welfare that maps to observable metrics, while also introducing popularity incentives as an unexplored channel in social networks distinct from the canonical mechanisms of conformity, learning, persuasion, and (mis)information transmission.
@@ 1  Introduction
Social media platforms are defined by a distinctive and universal feature: the quantification of social approval through “likes”. This mechanism is often tied to monetary or non-monetary rewards ([eckles2016estimating], [burtch2022peer], [zeng2023impact], [aridor2024economics], [filippas2025production]), systematically incentivizing agents for popularity—measured by likes received—of their content. When applied to opinion expression, this feature alters the strategic landscape: incentivizing agents to trade off popularity against authenticity.
@@
We develop a novel utilitarian framework for strategic expression that links utility to content production and exposure through observable social media metrics. We show that social media interaction acts as a state-dependent welfare amplifier. Moreover, strategic expression under popularity incentives systematically reshape the landscape of online discourse, generating redistributive and/or magnifying effects on individual utilities and aggregate welfare.
Our framework departs from existing models of content transmission—whether through network diffusion (e.g., [allcott2017social, vosoughi2018spread, steinert2015online], [abreu2019homophily]), algorithmic filtering (e.g., [levy2021social, guess2023social, bakshy2015exposure]), selective sharing (e.g., [campbell2019social, pogorelskiy2019news, garz2020partisan]), or spread of misinformation (e.g., [guriev2023curtailing, acemoglu2024model])—and instead models it as a popularity-driven environment.
In our model, agents hold heterogeneous and fixed authentic opinions, yet derive utility from social media activities: (i) utility gains from the popularity of their own posts, and (ii) utility gains (losses) from exposure to content that aligns with (diverges from) their authentic opinion. Abstracting from learning/belief updating [^2] Learning, belief updating, and persuasion have been widely studied in social networks literature (e.g., [degroot1974reaching, golub2010naive, golub2012network, jadbabaie2012non, molavi2018theory, grabisch2020survey]). Readers are referred to [pogorelskiy2019news] for a model of strategic information sharing in social media with learning in a voting framework. allows us to introduce a novel mechanism in social networks—strategic expression driven by popularity incentives under sticky opinions— and sheds light on a potential disconnect between societal beliefs and online discourse. Relatedly, some empirical studies show that social media often has limited effects on beliefs ([boxell2017greater, gentzkow2011ideological, nyhan2023like]).
@@
As emphasized in the review by [aridor2024economics], most of the existing literature evaluates welfare indirectly—through access choices, time use, or subjective well-being (e.g., [allcott2020welfare, allcott2022digital, brynjolfsson2023digital, braghieri2024political])—rather than specifying utility directly over observable metrics including self-posts, exposure to others’ posts, and engagement level for these posts. Our utility framework fills a critical gap in the literature333[aridor2024economics] emphasize in their review: …While most existing work has concentrated on the time spent on social media, the content consumed and the nature of the engagement are also important. Existing work has primarily quantified welfare through stated-preference measures, such as willingness to pay or accept for data sharing and privacy, which face well-known challenges including uncertainty, externalities, and gaps between stated and revealed preferences. Our framework directly addresses this gap while isolating from learning/persuasion channels. Developing a hybrid microfoundation that combines popularity-driven incentives with learning/persuasion is a natural direction for future research. by providing a micro-foundation for user welfare that maps directly to these observable platform metrics.
@@
To purely isolate the popularity-driven expression, we first focus on a representative social media environment in which each agent’s audience and exposure mirror the societal distribution of opinions. This benchmark abstracts from friendship and/or algorithmic biases—such as homophily—and clearly identifies the mechanism. In Section 6, we extend our analysis to algorithmic filtering by modeling a targeting rule widely used across social media platforms.
@@
We show a fundamental duality in how social media polarization differs from societal polarization: When society is already divided (e.g., political debates, immigration policy debates), popularity incentives amplify polarization in social media by inducing neutral agents to post opinionated content. Conversely, when society is highly unified (e.g., national celebrations, national sports events), the same incentives drive opinionated agents to converge toward neutral expression, generating amplified online unification. Social media thus acts as a non-linear magnifier, selectively distorting the perception of societal division depending on the nature of the event. Unlike existing work that studies online polarization (e.g., [gentzkow2011ideological, bakshy2015exposure, barbera2015tweeting, campbell2019social, pogorelskiy2019news, callander2022cause, guess2023reshares, della2023affective, vosoughi2018spread, acemoglu2024model, bolletta2025dynamic]), our results focuses on the role of popularity-driven strategic expression in it. Among these studies, [campbell2019social] analyze polarization in a social media network with two news content types—mass-market and niche—where users differ in their preferred content to recommend. In their model, individuals recommend their preferred type of content when it is available in their network, and otherwise recommend the alternative one. They show that when the friendship network is not biased, content filtering tends to amplify mass-market content and reduce polarization at steady state—–defined as the difference between the mass and niche content recommendation probabilities. However, greater network connectivity and homophily may increase the prevalence of niche content and, in turn, increase polarization. In another related work, [pogorelskiy2019news] analyzes content filtering in a similar two-type model within a voting framework. In their model, agents receive private signals and strategically share content based on their own partisan views and the composition of their audience on social media. Sharing decisions are strategic, whereas voting decisions are sincere and reflect Bayesian updating of beliefs about the state of the world. Different from these two models, our framework, in a multiple opinion setup, allows agents to strategically choose what to post (from a continuous opinion space) based on popularity and alignment incentives. Moreover, in a three-opinion setting, we show that the impact of strategic expression on polarization is state-dependent: polarization is amplified (dampened) when the underlying society is highly polarized (less polarized). Unlike [campbell2019social], network density plays no role in shaping polarization in our model, though it has sharp implications for user welfare, a dimension not addressed in these two studies.
@@
As emphasized by [aridor2024economics], an open question in the literature is estimating the effect of the intensity of social media usage on welfare. Our framework provides a state-dependent explanation with non-monotonicity, showing how different exposure levels can generate either amplified welfare gains or amplified welfare losses depending on topic intensity and the degree of polarization on the given topic.
@@
Our popularity trap complements the “participation trap” highlighted by [bursztyn2025product]. While their mechanism operates at the participation margin, ours arises at the margin of content creation within the platform. In their setting, agents feel compelled to join social media to avoid social exclusion, even if they derive negative utility from joining the platform that could be eliminated through coordinated exit. The popularity trap arising in our model could be avoided under coordinated authentic expression, suggesting some agents would be better off if they could collectively commit to authenticity in their posts.
Relatedly, our framework is distinct from the widely studied conformity models (e.g., [asch1955opinions, bernheim1994theory, morris2001political, zhang2018fashion, buechel2015opinion]). In conformity models, utility typically increases with similarity to others’ actions. These incentives generate coordination benefits: an agent benefits when others also conform to the same norm. In our framework, the direction of externalities is reversed. Agents do not desire collective convergence; rather, they prefer to enjoy higher utility from popularity, but prefer that other agents who share their authentic opinion continue to express it authentically. This generates distinct utility and welfare implications compared to conformity models, and also introduces the popularity traps.
@@ 3  Expressed Opinions and Distorted Utilities
Recall that the strategy space allows agents to post any opinion $c_{i}\in[-b,+b]$. However, Proposition 1 implies that popularity incentives do not generate novel viewpoints: no equilibrium post lies strictly between the discrete elements of the realized opinion set $\mathcal{O}$. Thus, despite the continuous choice set, expressed opinions are always constrained to the existing finite set of authentic beliefs. Popularity-driven social media, therefore, shapes outcomes not by creating new opinions, but by distorting the frequencies of existing ones.
Proposition 1 relies on a duality in equilibrium: agents’ authentic preferences are reflected in their likes on others’ posts but not necessarily in their own posts. At equilibrium, agent $i$ likes agent $j$’s post if and only if the post aligns with her own opinion, $c_{j}=b_{i}$. For any $c_{j}\neq b_{i}$, liking $j$’s post yields disutility $\omega^{d}_{i}|b_{i}-c_{j}|$, and therefore $i$ chooses not to react ($r_{ij}=0$). In contrast, when $j$’s post aligns with $i$’s opinion ($c_{j}=b_{i}$), liking provides a positive payoff $\omega^{a}_{i}$. Hence, $r_{ij}=1$ if and only if $c_{j}=b_{i}$.
@@
As a result, agent $i$’s equilibrium post $c_{i}$ solves:
$\max_{c_{i}\in[-b,+b]}A_{i,c_{i}}\bigl(\omega^{p}_{i}+\omega^{a}_{i}\mathbf{1}_{\{c_{i}=b_{i}\}}-\omega^{d}_{i}|b_{i}-c_{i}|\bigr),$
where $F_{i}$ captures agent $i$’s social media utility on the given topic except those derived from her own post.
@@
While the distortion mechanism applies to various distributions of societal opinions, in practice, societies often exhibit specific opinion distributions. For instance, during national celebrations or collective tragedies, a neutral opinion is held by a large majority, reflecting a highly unified society. In others, divisive events—ranging from high-stakes debates to low-stake daily-life or taste-based discussions on food, art, fashion,…—are characterized by polarization, with society mainly split across opposing positions. To capture these canonical environments and to characterize who benefits and who loses from social media interaction, we specialize to a three-opinion space in the remainder of the paper.
@@ 5  Individual Utilities and Welfare
We now systematically compare the utility difference between the PE and the authentic benchmark. Before showing formal existence-based results, first we provide the full characterization of potential regimes.
@@ 6  An Algorithmic Tool: Preference-Based Exposure
Next, we revisit the timing of the model. Posts are created at time $t=0$, and at this step, agents are assumed to know the platform’s algorithmic rule [^12] While the assumption that agents know the algorithm may appear strong, it is consistent with the mechanics of like-based suggestions widely used by social media platforms ([bakshy2015exposure]). This feedback—where likes reveal users’ preferences and the platform adjusts visibility in response—illustrates how platforms gradually learn user types. On platforms that use like-based suggestions, it is therefore reasonable for agents to expect the algorithm to behave like a post-viewer-match rule, showing posts primarily to users whose opinions align with the content.. After posts are created, the platform chooses the set $\mathcal{V}_{i}$ for each $i\in\mathcal{N}$ at time $t=1$, before making these posts visible on social media at time $t=2$. The information structure and the timing of posting actions and reactions to posts are same as in Section 2.
@@
Given post $c_{i}$, the platform chooses the set $\mathcal{V}_{i}$ for each $i$ that maximizes the number of likes subject to the visibility constraint $V_{i}\leq V^{\star}_{i}$ (referring to engagement maximization by platform):
$\max_{\mathcal{V}_{i}\subseteq\mathcal{N}}\;\sum_{j\in\mathcal{V}_{i}}r_{ji}\quad\text{s.t.}\quad V_{i}\leq V^{\star}_{i}.$
@@
Representative algorithm (RA). This algorithm exactly matches the representative exposure, but under a visibility cap. Under this algorithm, each agent $i$’s post is shown to $V_{i}\cdot\frac{G_{b^{\star}}}{n}$ number of agents from each opinion group $b^{\star}\in\mathcal{O}$. [^13] We assume that parameters in $V_{i}\cdot\frac{G_{b^{\star}}}{n}$ satisfy conditions such that there is no divisibility issue. Consequently, the total likes of agent $i$ is $R_{i}=V_{i}\cdot\frac{G_{c_{i}}}{n}$, where $G_{c_{i}}$ is the group size of agents in society holding the view that is same as $c_{i}$. This implies the platform maximizes the number of likes by setting $V_{i}=V^{\star}_{i}$, and hence, total likes for each agent $i$ is:
$R_{i}=V^{\star}_{i}\cdot\frac{G_{c_{i}}}{n}.$
@@
  • Suppose the baseline utility, utility parameters, and visibility caps are homogeneous: $H_{i}=\overline{H},\quad\omega_{i}^{p}=\omega^{p},\quad\omega_{i}^{a}=\omega^{a},\quad\omega_{i}^{d}=\omega^{d},\quad a_{i}=a,\quad V_{i}^{*}=k\text{ }\forall i\in\mathcal{N}.$
  • i)

### v2

@@ Abstract
Social media platforms systematically reward popularity over authenticity, incentivizing users to strategically tailor their expression for attention. In this paper, we introduce (i) an attention-seeking model, distinct from canonical mechanisms of conformity, learning, persuasion, and (mis)information transmission in social networks literature, and (ii) the first utilitarian framework defined directly over observable social media platform metrics, filling a critical gap in the social media literature. In the model, agents hold fixed heterogeneous authentic opinions and derive (i) utility gains from the popularity of their own posts—measured by likes received, and (ii) utility gains (losses) from exposure to content that aligns with (diverges from) their authentic opinion. Social media interaction acts as a state-dependent welfare amplifier: light topics generate Pareto improvements, whereas intense topics make everyone worse off in a polarized society (e.g., political debates during elections). Moreover, strategic expression amplifies social media polarization during polarized events while dampening it during unified events (e.g., national celebrations). Consequently, strategic distortions magnify welfare outcomes, expanding aggregate gains in light topics while exacerbating losses in intense, polarized ones. Counterintuitively, strategic agents often face a popularity trap: posting a more popular opinion is individually optimal, yet collective action by similar agents eliminates their authentic opinion from the platform, leaving them worse off than under the authentic-expression benchmark. Preference-based algorithms—widely used by platforms— or homophilic exposures discipline popularity-driven behavior, narrowing the popularity trap region and limiting its welfare effects.
@@ 1  Introduction
Social media platforms are defined by a distinctive feature: the quantification of social approval through “likes”. This mechanism is often tied to monetary or non-monetary rewards ([23], [18], [42], [6], [24]), thereby incentivizing users to seek popularity for their content. When applied to opinion expression, this feature alters the strategic landscape: incentivizing agents to trade off popularity against authenticity.
@@
In this paper, we study strategic expression in social media. We show that social media acts as a state-dependent welfare amplifier. Moreover, popularity-driven strategic expression systematically reshapes the landscape of online discourse, generating redistributive and/or magnifying effects on polarization, individual utilities, and aggregate welfare.
Our main contribution is twofold. First, our model introduces popularity (or attention-seeking) as an unexplored channel in the social networks literature, distinct from the canonical mechanisms of conformity, learning, belief updating, and persuasion (e.g., [7, 21, 10, 36, 25, 28, 33, 35, 11, 29]) [^2] Readers are referred to [39] for a model of strategic information sharing in social media with learning in a voting framework.. In our analysis, we abstract away from these canonical mechanisms to discover this new strategic landscape, shedding light on a potential disconnect between beliefs and online discourse. Relatedly, empirical studies show that social media often has limited effects on beliefs ([27, 13, 38]).
@@
Second, we are the first to introduce a utilitarian framework defined directly over social media metrics: posts and likes. As emphasized in the review by [6], the existing literature evaluates welfare indirectly—through access choices, time use, or subjective well-being (e.g., [3, 4, 15, 14])—rather than specifying utility directly over observable metrics. Our tractable utilitarian framework fills this critical gap in the social media literature. This workhorse model allows us to measure individual utilities and aggregate welfare directly from social media interaction.333[6] emphasize in their review: …While most existing work has concentrated on the time spent on social media, the content consumed and the nature of the engagement are also important. Existing work has primarily quantified welfare through stated-preference measures, such as willingness to pay or accept for data sharing and privacy, which face well-known challenges including uncertainty, externalities, and gaps between stated and revealed preferences. Our framework directly addresses this gap while isolating from learning/persuasion channels. Developing a hybrid microfoundation that combines popularity-driven incentives with learning/persuasion is a natural direction for future research.
@@
In our model, agents hold heterogeneous and fixed authentic opinions, yet derive utility from social media activities: (i) utility gains from the popularity of their own posts, and (ii) utility gains (losses) from exposure to content that aligns with (diverges from) their authentic opinion. To purely isolate the popularity-driven expression, we first focus on a representative social media environment in which each agent’s audience and exposure mirror the societal distribution of opinions. This benchmark abstracts from friendship and/or algorithmic biases—such as homophily—and clearly identifies the mechanism. In Section 6, we extend our analysis to algorithmic filtering by modeling a targeting rule widely used across social media platforms.
@@
We show a fundamental duality in how social media polarization differs from societal polarization: When society is already divided (e.g., political debates, immigration policy debates), popularity incentives amplify polarization in social media by inducing neutral agents to post opinionated content. Conversely, when society is highly unified (e.g., national celebrations, national sports events), the same incentives drive opinionated agents to converge toward neutral expression, generating amplified online unification. Social media thus acts as a non-linear magnifier, selectively distorting the perception of societal division depending on the nature of the event. Unlike existing work that studies online polarization (e.g., [27, 8, 9, 20, 39, 19, 31, 22, 41, 2, 12]), our results focus on the role of popularity-driven strategic expression in it. Among these studies, [20] analyze polarization in a social media network with two news content types—mass-market and niche—where users differ in their preferred content to recommend. In their model, individuals recommend their preferred type of content when it is available in their network, and otherwise recommend the alternative one. They show that when the friendship network is not biased, content filtering tends to amplify mass-market content and reduce polarization at steady state— defined as the difference between the mass and niche content recommendation probabilities. However, greater network connectivity and homophily may increase the prevalence of niche content and, in turn, increase polarization. In another related work, [39] analyzes content filtering in a similar two-type model within a voting framework. In their model, agents receive private signals and strategically share content based on their own partisan views and the composition of their audience on social media. Sharing decisions are strategic, whereas voting decisions are sincere and reflect Bayesian updating of beliefs about the state of the world. Different from these two models, our framework, in a multiple opinion setup, allows agents to strategically choose what to post (from a continuous opinion space) based on popularity and alignment incentives. Moreover, in a three-opinion setting, we show that the impact of strategic expression on polarization is state-dependent: polarization is amplified (dampened) when the underlying society is highly polarized (less polarized). Unlike [20], network density plays no role in shaping polarization in our model, though it has sharp implications for user welfare, a dimension not addressed in these two studies.
@@
As emphasized by [6], an open question in the literature is estimating the effect of the intensity of social media usage on welfare. Our framework provides a state-dependent explanation with non-monotonicity, showing how different exposure levels can generate either amplified welfare gains or amplified welfare losses depending on topic intensity and the degree of polarization on the given topic.
@@
Our popularity trap complements the “participation trap” highlighted by [17]. While their mechanism operates at the participation margin, ours arises at the margin of content creation within the platform. In their setting, agents feel compelled to join social media to avoid social exclusion, even if they derive negative utility from joining the platform that could be eliminated through coordinated exit. The popularity trap arising in our model could be avoided under coordinated authentic expression, suggesting some agents would be better off if they could collectively commit to authenticity in their posts.
Relatedly, our framework is distinct from the widely studied conformity models (e.g., [7, 10, 36, 43, 16]). In conformity models, utility typically increases with similarity to others’ actions. These incentives generate coordination benefits: an agent benefits when others also conform to the same norm. In our framework, the direction of externalities is reversed. Agents do not desire collective convergence; rather, they prefer to enjoy higher utility from popularity, but prefer that other agents who share their authentic opinion continue to express it authentically. This generates distinct utility and welfare implications compared to conformity models, and also introduces the popularity traps.
Finally, our framework departs from existing models of content transmission online—whether through network diffusion (e.g., [5, 41, 40], [1]), algorithmic filtering (e.g., [34, 30, 8]), selective sharing (e.g., [20, 39, 26]), or spread of misinformation (e.g., [37, 32, 2])—and instead models it as a popularity-driven environment.
@@ 3  Expressed Opinions and Distorted Utilities
Recall that the strategy space allows agents to post any opinion $c_{i}\in[-b,+b]$. However, Lemma 1 implies that popularity incentives do not generate novel viewpoints: no equilibrium post lies strictly between the discrete elements of the realized opinion set $\mathcal{O}$. Thus, despite the continuous choice set, expressed opinions are always constrained to the existing finite set of authentic beliefs. Popularity-driven social media, therefore, shapes outcomes not by creating new opinions, but by distorting the frequencies of existing ones.
Lemma 1999The formal proof of Lemma 1 and the proof of Proposition 6 are in the Online Appendix. relies on a duality in equilibrium: agents’ authentic preferences are reflected in their likes on others’ posts but not necessarily in their own posts. At equilibrium, agent $i$ likes agent $j$’s post if and only if the post aligns with her own opinion, $c_{j}=b_{i}$. For any $c_{j}\neq b_{i}$, liking $j$’s post yields disutility $\omega^{d}_{i}|b_{i}-c_{j}|$, and therefore $i$ chooses not to react ($r_{ij}=0$). In contrast, when $j$’s post aligns with $i$’s opinion ($c_{j}=b_{i}$), liking provides a positive payoff $\omega^{a}_{i}$. Hence, $r_{ij}=1$ if and only if $c_{j}=b_{i}$.
@@
The baseline utility, $H_{i}$, and the utility component $F_{i}$—that captures agent $i$’s social media utility on the given topic except those derived from her own post— are both independent from the agent’s own posting decision, and therefore do not play a role in the posting decision of agent $i$. As a result, agent $i$’s equilibrium post $c_{i}$ solves:
$\max_{c_{i}\in[-b,+b]}A_{i,c_{i}}\bigl(\omega^{p}_{i}+\omega^{a}_{i}\mathbf{1}_{\{c_{i}=b_{i}\}}-\omega^{d}_{i}|b_{i}-c_{i}|\bigr).$
@@
While the distortion mechanism applies to various distributions of societal opinions, in practice, societies often exhibit specific opinion distributions. For instance, during national celebrations or collective tragedies, a neutral opinion is held by a large majority, reflecting a highly unified society. In others, divisive events—ranging from high-stakes debates to low-stake daily-life or taste-based discussionson on food, art, fashion,…—are characterized by polarization, with society mainly split across opposing positions. To capture these canonical environments and to characterize who benefits and who loses from social media interaction, we specialize to a three-opinion space in the remainder of the paper.
@@ 5  Individual Utilities and Welfare
We now systematically compare the utility difference between the popularity-driven equilibrium and the authentic benchmark. Before showing formal existence-based results, first we provide the full characterization of potential regimes.
@@ 6  An Algorithmic Tool: Preference-Based Exposure
Next, we revisit the timing of the model. Posts are created at time $t=0$, and at this step, agents are assumed to know the platform’s algorithmic rule [^13] While the assumption that agents know the algorithm may appear strong, it is consistent with the mechanics of like-based suggestions widely used by social media platforms ([8]). This feedback—where likes reveal users’ preferences and the platform adjusts visibility in response—illustrates how platforms gradually learn user types. On platforms that use like-based suggestions, it is therefore reasonable for agents to expect the algorithm to behave like a post-viewer-match rule, showing posts primarily to users whose opinions align with the content.. After posts are created, the platform chooses the set $\mathcal{V}_{i}$ for each $i\in\mathcal{N}$ at time $t=1$, before making these posts visible on social media at time $t=2$. The information structure and the timing of posting actions and reactions to posts are same as in Section 2.
@@
Given post $c_{i}$, the platform chooses the set $\mathcal{V}_{i}$ for each $i$ that maximizes the number of likes subject to the visibility constraint $V_{i}\leq V^{\ast}_{i}$ (referring to engagement maximization by platform):
$\max_{\mathcal{V}_{i}\subseteq\mathcal{N}}\;\sum_{j\in\mathcal{V}_{i}}r_{ji}\quad\text{s.t.}\quad V_{i}\leq V^{\ast}_{i}.$
@@
Representative algorithm (RA). This algorithm exactly matches the representative exposure, but under a visibility cap. Under this algorithm, each agent $i$’s post is shown to $V_{i}\cdot\frac{G_{b^{\ast}}}{n}$ number of agents from each opinion group $b^{\ast}\in\mathcal{O}$. [^14] We assume that parameters in $V_{i}\cdot\frac{G_{b^{\ast}}}{n}$ satisfy conditions such that there is no divisibility issue. Consequently, the total likes of agent $i$ is $R_{i}=V_{i}\cdot\frac{G_{c_{i}}}{n}$, where $G_{c_{i}}$ is the group size of agents in society holding the view that is same as $c_{i}$. This implies the platform maximizes the number of likes by setting $V_{i}=V^{\ast}_{i}$, and hence, total likes for each agent $i$ is:
$R_{i}=V^{\ast}_{i}\cdot\frac{G_{c_{i}}}{n}.$
@@
  • Suppose the baseline utility, utility parameters, and visibility caps are homogeneous: $H_{i}=\overline{H},\quad\omega_{i}^{p}=\omega^{p},\quad\omega_{i}^{a}=\omega^{a},\quad\omega_{i}^{d}=\omega^{d},\quad a_{i}=a,\quad V_{i}^{*}=k\quad\forall i\in\mathcal{N}.$
  • (i)